The average yearly earnings of engineering majors exceed the earnings of marketing majors by $24,000. The average yearly earnings of accounting majors exceed the earnings of marketing majors by $5,000. Combined, the average yearly earnings for these three college majors are $245,000. Determine the average yearly earnings, in thousands of dollars, for each of these three college majors.
Answer by greenestamps(13198)



The given information compares the average salaries of both engineering majors and accounting majors to the average of marketing majors, so it makes sense to use the average salary of marketing majors as the "baseline".

x = average salary of marketing majors (thousands of dollars)
x+5 = average salary of accounting majors
x+24=average salary of engineering majors

x+(x+5)+(x+24)=245
3x+29=245
3x=216
x=72

ANSWERS:
average salary for marketing majors: x=$72,000
average salary for accounting majors: x+5=$77,000
average salary for accounting majors: x+24=$96,000